# Multimodal Crash Likelihood Prediction: A Complexity-Infused Approach Integrating Semantic, Contextual, and Driving Features

### v1

@@ 4.1 Complexity-infused Encoder
Implementation. The complexity-infused encoder was implemented with different settings, depending on the input feature sets and the type of output variable, i.e., the complexity index. When the complexity index was treated as a continuous variable, the Mean Squared Error (MSE) was used as the loss function. When the index was treated as a categorical variable, Cross Entropy (CE) was chosen as the loss function. Further details of the model training configurations can be found in the Appendix.
Results. The model performance of the encoder is presented in Table 1, showing the model performance on the complexity index obtained from the LLMs. As shown in the table, the three model settings exhibited similar performance across different input feature sets. However, several insights can be drawn from the results: first, the model with 32 hidden neurons and a continuous output produced the lowest RMSE overall. Second, models with more comprehensive input feature sets, which incorporated semantic, driving, and contextual features, consistently outperformed those with fewer feature sets. This pattern was observed across all three models.
@@
Semantic + Driving $1.05$ $1.06$
Sem. + Driv. + Cont. $0.84$ $0.86$
@@
Semantic + Driving $1.08$ $1.08$
Sem. + Driv. + Cont. $0.84$ $0.85$
@@
Semantic + Driving $1.31$ $1.34$
Sem. + Driv. + Cont. 1.15.15.15.15 $1.18$
Table 1: Performance comparison of the complexity-infused encoder models with different input feature sets and hidden neurons. The results are reported in terms of RMSE for continuous outputs and Cross-Entropy for categorical outputs.
Since the encoder with 32 hidden neurons and a continuous output variable yielded the best results, the 32 hidden features from this model will be used as the complexity-infused features in Section 4.2.
@@ 4.2 Main Results of Crash Prediction Model
Implementation. When training the neural network for the crash prediction model, Cross Entropy was chosen as the loss function during training. The accuracy of the best-performing model will be reported in the subsequent section. Further details on the model training configurations can be found in the Appendix.
@@
Semantic + Driving + Contextual RF 87.98 90.15 2.17 78.49 87.35
@@
The results indicated that the Random Forest model consistently achieved the best performance across all combinations of input feature sets. Additionally, there was a clear trend of improved model performance as the number of input features increased, with the highest accuracy reaching 87.98%. Adding the complexity-infused features led to further improvements in prediction performance, with the highest accuracy being 90.15%, when using semantic + driving + contextual features. This was followed by the neural network model, which achieved an accuracy of $88.78\%$. To assess the significance of these improvements, McNemar’s test [14] was conducted for each model, and the results indicated that the improvements were statistically significant. The results suggested that incorporating complexity-infused features, particularly in models using a combination of semantic, driving, and contextual features, significantly enhances the accuracy of crash likelihood predictions.
For the best-performing model, which was trained on all available features, including complexity-infused features, using the Random Forest algorithm, the feature importance is shown in Fig. 6. The results indicate that, among the semantic features, the most important factors in classifying crash likelihood include the vegetation percentage, sky percentage, road percentage, car count, and car percentage, both in the full frame and lead-car region.
For driving features, key factors include the original speed, mean and standard deviation of speed, raw and normalized speed deviation from the speed limit, and the standard deviation of longitudinal acceleration. Additionally, four hidden complexity-infused features (labeled 0, 1, 21, and 30) were identified as important. Notably, none of the contextual features were ranked among the top 20 most important features in the crash prediction model.
Figure 6: The feature importance of the best-performing model.
Using the complexity-infused features alone or adding the 1-dimensional complexity index to the baseline models did not outperform the baseline models, except for the neural network model on Semantic + Driving + Contextual features. This suggests that simply adding the complexity index does not substantially improve the model’s prediction capability. However, the complexity-infused features were found to effectively enhance model performance. This indicates that the complexity index is closely associated with scene-level semantic features, driving behavior, and contextual characteristics, and that hidden information can be effectively extracted using neural network structures.
@@ 4.3 Ablation Studies
4.3.1 Comparison of LLMs and MTurk Annotations
The encoder trained on the MTurk annotations utilized a 32-hidden neuron structure and treated the output as a continuous variable, since the annotations were generated by averaging the workers’ responses. The performance of the encoders trained on LLMs and MTurk data is shown in Table 3. It can be seen that the encoder using MTurk data performed slightly better than the one using LLMs data. However, for the encoders trained on the combination of semantic, driving, and contextual features, the performance of both encoders was very similar.
@@ 4.3.1 Comparison of LLMs and MTurk Annotations
LLMs Semantic 1.10 1.10
Semantic+Driving 1.05 1.06
Sem. + Driv. + Cont. 0.84 0.86
@@
Semantic+Driving 0.91 0.93
Sem. + Driv. + Cont. 0.83 0.88
Table 3: Performance comparison of the complexity-infused encoder models using LLMs data or MTurk data. The results are reported in terms of RMSE. Downward arrows next to the values indicate that lower RMSE is preferable.
Next, the crash prediction model was trained using the complexity-infused features derived from the MTurk data. The results of the crash likelihood prediction models trained on both LLMs and MTurk data are shown in Table 4. The model using complexity-infused features derived from LLMs data consistently yielded higher accuracy than the one using MTurk data. This suggests that the LLMs-generated complexity indexes may capture relevant information more effectively for crash likelihood prediction.
@@
LLMs Semantic 78.35 65.31
Semantic+Driving 86.32 67.85
Sem. + Driv. + Cont. 90.15 78.49
@@
Semantic+Driving 80.89 73.61
Sem. + Driv. + Cont. 86.60 80.06
Table 4: Performance comparison of the crash prediction models using complexity-infused features from LLMs data or MTurk data. The results are reported in terms of accuracy (%). Upward arrows next to the values indicate that higher accuracy is preferable.
It is also worth noting that the model using only the complexity-infused features from MTurk annotations outperformed the one based on LLMs annotations. When considering the encoder performance (Table 3), the model trained solely on complexity-infused features from MTurk data showed better prediction capability than the ones from LLMs data. However, when semantic, driving, and contextual features were added, the best-performing model was achieved using the LLMs data. This implies that while the complexity-infused features derived from MTurk annotations capture important information effectively in isolation, the LLMs-generated data provides a more robust integration of semantic, driving, and contextual features. As a result, the combination of LLMs data with these additional features enhances the model’s overall performance.
@@ 4.3.2 Comparisons of Other Encoders
Semantic + Driving 86.32
Sem. + Driv. + Cont. 90.15
@@
Semantic + Driving 86.24
Sem. + Driv. + Cont. 89.56
@@
Semantic + Driving 83.59
Sem. + Driv. + Cont. 88.65
@@
The models using different encoders consistently achieved higher accuracy as the number of input features increased. The encoder with 32 hidden neurons and a continuous output consistently delivered the highest accuracy, regardless of the input feature set. These results further validate the proposed encoder structure’s effectiveness in extracting hidden context from the existing semantic, driving, and contextual features through the complexity index.
@@ 5 Conclusions and Limitations
In this paper, we presented a two-stage framework for extracting hidden context from semantic, driving, and contextual features, and for predicting crash likelihood by incorporating these hidden context features into the original feature sets. This approach addresses the challenges of predicting crashes in real-world and naturalistic settings. To our knowledge, this is the first model to integrate all scene-related and driving-related features and link them to a complexity index obtained from the LLMs model. Our experiments revealed that the framework can accurately predict crash likelihood, achieving an accuracy of $90.15\%$. The data generated by LLMs provided better predictive capability than that obtained from MTurk, and different encoder structures consistently outperformed models without complexity-infused features. This solution has the potential to inform the development of warning systems and driving monitoring systems in human-machine interfaces, improve safety in autonomous driving, and enhance human-AI collaboration.
@@
In the future, we aim to expand the dataset by incorporating more diverse driving environments and conditions, such as adverse weather and complex urban settings, to further improve the generalizability of the model. In addition, we intend to experiment with more advanced encoder architectures to further refine the extraction of hidden context features. Finally, we plan to investigate the integration of real-time data, enabling dynamic crash likelihood predictions for real-world applications.

### v2

@@ 4 Experiments and Results
This section presents the modeling results of the encoder and the crash density prediction models. Section 4.1 and Section 4.2 discuss the results obtained using AI-generated annotations, specifically from the GPT model, as the crash density prediction models utilizing AI annotations consistently outperformed those using human annotations. Section 4.3 provides a comparative analysis of AI annotations versus human annotations, as well as an evaluation of different encoder structures.
@@ 4.1 Complexity-infused Encoder
Implementation. The complexity-infused encoder was implemented with different settings, depending on the input feature sets and the type of output variable, i.e., the complexity index. When the complexity index was treated as a continuous variable, the Root Mean Squared Error (RMSE) was used as the loss function. When the index was treated as a categorical variable, Cross Entropy (CE) was chosen as the loss function. Further details of the model training configurations can be found in Appendix 10.1.
Results. The model performance of the encoder is presented in Table 1, showing the model performance on the complexity index obtained from the LLM. As shown in the table, the three model settings exhibited similar performance across different input feature sets. However, several insights can be drawn from the results: first, the model with 32 hidden neurons and a continuous output produced the lowest RMSE overall. Second, models with more comprehensive input feature sets, which incorporated semantic, kinematic, and contextual features, consistently outperformed those with fewer feature sets. This pattern was observed across all three models.
@@
Semantic + kinematic $1.05$ $1.06$
Sem. + Kin. + Cont. $0.84$ $0.86$
@@
Semantic + kinematic $1.08$ $1.08$
Sem. + Kin. + Cont. $0.84$ $0.85$
@@
Semantic + kinematic $1.31$ $1.34$
Sem. + Kin. + Cont. 1.15.15.15.15 $1.18$
Table 1: Performance comparison of the complexity-infused encoder models with different input feature sets and hidden neurons. The results are reported in terms of RMSE for continuous outputs and Cross-Entropy for categorical outputs. Downward arrows next to the values indicate that lower values are preferable.
Since the encoder with 32 hidden neurons and a continuous output variable yielded the best results, the 32 hidden features from this model were used as the complexity-infused features in Section 4.2.
@@ 4.2 Main Results of Crash Prediction Model
Implementation. When training the neural network for the crash prediction model, Cross Entropy was chosen as the loss function during training. The accuracy of the best-performing model is reported in the next section. Further details on the model training configurations can be found in Appendix 10.2.
@@
Semantic + kinematic + Contextual RF 87.98 90.15 2.17 78.49 87.35
@@
Figure 6: The SHAP values of the 10 most influential features for each class of the best-performing model.
The results indicated that the Random Forest model consistently achieved the best performance across all combinations of input feature sets. Additionally, there was a clear trend of improved model performance as the number of input features increased, with the highest accuracy reaching 87.98%. Adding the complexity-infused features led to further improvements in prediction performance, with the highest accuracy being 90.15%, when using semantic, kinematic, and contextual features. This was followed by the neural network model, which achieved an accuracy of $88.78\%$. To assess the significance of these improvements, McNemar’s test [16] was conducted for each model, and the results indicated that the improvements were statistically significant. The results suggested that incorporating complexity-infused features, particularly in models using a combination of semantic, kinematic, and contextual features, significantly enhances the accuracy of crash likelihood predictions.
For the best-performing model, which was trained on all available features, including complexity-infused features, using the Random Forest algorithm, the SHAP values of the 10 most influential features for each class are shown in Fig. 6. The SHAP values highlight the features that most strongly differentiate between each pair of classes. For example, when differentiating the low-density class from the medium-density class, the vegetation area in both the full frame and the lead-car region emerged as the most significant features. Higher values of vegetation area were associated with low crash density areas.
For the high-density class, speed-related features were the most influential. The results revealed that lower speeds are associated with higher crash-density areas. Regarding contextual features, rural roads were associated with medium crash-density areas compared to low-density areas, while highways were more frequently associated with high crash-density areas. Among the complexity-infused features, at least one was found to be influential in differentiating between low and high crash-density areas.
Using the complexity-infused features alone or adding the 1-dimensional complexity index to the baseline models did not outperform the baseline models, except for the neural network model on semantic, kinematic, and contextual features. This suggests that simply adding the complexity index does not substantially improve the model’s prediction capability. However, the complexity-infused features were found to effectively enhance model performance. This indicates that the complexity index is closely associated with semantic representations, vehicle kinematics, and contextual characteristics, and that hidden information can be effectively extracted using neural network structures.
@@ 4.3 Ablation Studies
4.3.1 Comparison of LLM and MTurk Annotations
The encoder trained on the MTurk annotations utilized a 32-hidden neuron structure and treated the output as a continuous variable, since the annotations were generated by averaging the workers’ responses. The performance of the encoders trained on LLM and MTurk data is shown in Table 3. It can be seen that the encoder using MTurk data performed slightly better than the one using LLM data. However, for the encoders trained on the combination of semantic, kinematic, and contextual features, the performance of both encoders was very similar.
@@ 4.3.1 Comparison of LLM and MTurk Annotations
LLM Semantic 1.10 1.10
Semantic + kinematic 1.05 1.06
Sem. + Kin. + Cont. 0.84 0.86
@@
Semantic + kinematic 0.91 0.93
Sem. + Kin. + Cont. 0.83 0.88
Table 3: Performance comparison of the complexity-infused encoder models using LLM data or MTurk data. The results are reported in terms of RMSE. Downward arrows next to the values indicate that lower RMSE is preferable.
Next, the crash prediction model was trained using the complexity-infused features derived from the MTurk data. The results of the crash likelihood prediction models trained on both LLM and MTurk data are shown in Table 4. The model using complexity-infused features derived from LLM data consistently yielded higher accuracy than the one using MTurk data. This suggests that the LLM-generated complexity indexes may capture relevant information more effectively for crash likelihood prediction.
@@
LLM Semantic 78.35 65.31
Semantic + kinematic 86.32 67.85
Sem. + Kin. + Cont. 90.15 78.49
@@
Semantic + kinematic 80.89 73.61
Sem. + Kin. + Cont. 86.60 80.06
Table 4: Performance comparison of the crash prediction models using complexity-infused features from LLM data or MTurk data. The results are reported in terms of accuracy (%). Upward arrows next to the values indicate that higher accuracy is preferable.
It is also worth noting that the model using only the complexity-infused features from MTurk annotations outperformed the one based on LLM annotations. When considering the encoder performance (Table 3), the model trained solely on complexity-infused features from MTurk data showed better prediction capability than the ones from LLM data. However, when semantic, kinematic, and contextual features were added, the best-performing model was achieved using the LLM data. This implies that while the complexity-infused features derived from MTurk annotations capture important information effectively in isolation, the LLM-generated data provides a more robust integration of semantic, kinematic, and contextual features. As a result, the combination of LLM data with these additional features enhances the model’s overall performance.
@@ 4.3.2 Comparisons of Other Encoders
Semantic + Kinematic 86.32
Sem. + Kin. + Cont. 90.15
@@
Semantic + Kinematic 86.24
Sem. + Kin. + Cont. 89.56
@@
Semantic + Kinematic 83.59
Sem. + Kin. + Cont. 88.65
@@
The models using different encoders consistently achieved higher accuracy as the number of input features increased. The encoder with 32 hidden neurons and a continuous output consistently delivered the highest accuracy, regardless of the input feature set. These results further validate the proposed encoder structure’s effectiveness in extracting hidden context from the existing semantic, kinematic, and contextual features through the complexity index.
@@ 5 Conclusions and Limitations
In this paper, we presented a two-stage framework for extracting hidden context from semantic, kinematic, and contextual features, and for predicting crash density by incorporating these hidden context features into the original feature sets. This approach addresses the challenges of identifying and understanding the key factors associated with crash prevalence in real-world naturalistic driving environments. To our knowledge, this is the first model to integrate all scene-related and driving-related features and link them to a complexity index obtained from the LLM. Our experiments revealed that the framework can accurately predict crash density, achieving an accuracy of $90.15\%$. The data generated by LLM provided better predictive capability than that obtained from human annotation via MTurk. Different encoder structures consistently outperformed models without complexity-infused features. This solution has the potential to inform the development of advanced driver assistance systems, and driver monitoring systems to improve safety in manually driven and automated vehicles. The work can lead to enhanced human-AI collaboration that more effectively supports drivers through complex environmental changes. Furthermore, insights from this framework have the potential to support roadway design by empowering highway engineering departments with the data needed to identify and mitigate risk factors in crash hotspots. If embraced through infrastructure changes these results would thus contribute to a key proactive traffic safety intervention that aligns with US DOT’s Safe-Systems Approach.
@@
In the future, to further improve the generalizability of the model, we aim to expand the dataset by incorporating more diverse driving environments and a broader range of driving conditions, such as adverse weather and complex urban settings. Furthermore, we intend to experiment with more advanced encoder architectures to refine the extraction of hidden context features. Finally, we plan to investigate the integration of real-time data, enabling dynamic crash likelihood predictions for real-world applications.
6 Acknowledgment
Data for this study were drawn from work supported by the Advanced Vehicle Technologies (AVT) Consortium at MIT (http://agelab.mit.edu/avt). The views and conclusions expressed are those of the authors and have not been sponsored, approved, or endorsed by the Consortium or Foundation.
This paper is funded partially by the New England University Transportation Center (NEUTC). NEUTC is funded by a grant from the U.S. Department of Transportation’s University Transportation Centers Program (69A3552348301). However, the U.S. Government assumes no liability for the contents or use thereof.
Supplementary Material
7 LLM Implementation Details
The contextual features and the complexity index feature mentioned in Sec. 3.2.3 and Sec. 3.3.1 were generated using the GPT-4o-2024-08-06 model  [^4] https://platform.openai.com/docs/models/gpt-4o. Batch API  [^5] https://platform.openai.com/docs/guides/batch was used for faster processing.
8 Annotation Comparisons
Table 6 presents the descriptive statistics for the comparison between AI-generated and human-generated annotated complexity indices, as discussed in Sec. 3.3.1. This comparison is based on the entire dataset without matching. It can be observed that the overall distribution of the complexity index annotated by AI and humans is notably similar.
Method M (SD) Min 25% 50% 75% Max
AI 4.2 (1.3) 1 3 4 5 8
Human 4.5 (1.2) 1.7 3.6 4.5 5.2 8.2
Table 6: The descriptive statistics of the complexity index from AI-generated and human-generated annotations.
When looking at the difference in complexity indices between AI and human annotations (diff $=$ complexity index of AI -- complexity index of human), the difference has a mean of $-0.25$, a standard deviation of $1.07$, a minimum of $-4.5$, a 25th percentile of $-1$, a median of $-0.3$, a 75th percentile of $0.4$, and a maximum of $3.2$. It can be observed that the differences are centered around zero with AI annotations being slightly smaller than human annotations on average.
9 Model Structure
The neural network model structure mentioned in Sec. 3.4.2 is shown in Fig. 7.
Input LayerFully Connected (256 units) + ReLU ActivationFully Connected (512 units) + ReLU ActivationDropout (50%)Fully Connected (256 units) + ReLU ActivationFully Connected (128 units) + ReLU ActivationFully Connected (64 units) + ReLU ActivationFully Connected (32 units) + ReLU ActivationFully Connected, Output UnitsSoftmax Output
Figure 7: Visualization of the crash density prediction neural network model architecture.
10 Model Training Details
10.1 Encoder Training Details
Regarding the training configurations of the encoders mentioned in Sec. 4.1, different setups were applied based on the input features and the output data types. The model was trained for $1000$ epochs using Stochastic Gradient Descent (SGD) as the optimizer with a momentum value of $0.9$. The learning rate varied depending on the feature and data type:
  • Semantic features only and continuous data type: Learning rate of $0.0005$.
  • Semantic features only and categorical data type: Learning rate of $0.005$.
  • Semantic and driving features and continuous data type: Learning rate of $0.0003$.
  • Semantic and driving features and categorical data type: Learning rate of $0.001$.
  • All available features (semantic, driving, and contextual) and continuous data type: Learning rate of $0.0005$.
  • All available features (semantic, driving, and contextual) and categorical data type: Learning rate of $0.001$.
A Cosine Annealing learning rate scheduler was used to adjust the learning rate over the training epochs.
10.2 Prediction Model Training Details
Regarding the training configurations of the prediction models mentioned in Sec. 4.2, all models were trained using SGD with a momentum value of $0.9$. The learning rate was set to $0.001$, with a Cosine Annealing scheduler. The total number of epochs was $2000$.